1) Use equilibrium ion concentration to calculate Ksp. The Pb2+ concentration in a saturated solution of...

1)

Use equilibrium ion concentration to calculate Ksp.

The Pb2+ concentration in a saturated solution of lead bromide is measured and found to be 1.19×10-2 M. Use this information to calculate a Ksp value for lead bromide.

Ksp =___________

2)

Use solubility to calculate Ksp.

The solubility of Fe(OH)2 is measured and found to be 1.15×10-3 g/L. Use this information to calculate a Ksp value for iron(II) hydroxide.

Ksp =______________
Answer #1

1) Lead bromide dissolves in water to a small extent according to the equation

PbBr2 = Pb^2+ + 2Br-

The concentration of Pb2+ = 1.19*10^-2M

if the concentration of Pb2+ is x, then concentration of Br- is 2x,i.e., 2.38*10^-2M

therefore, Ksp = (Pb2+)(Br-)^2 = (1.19*10^-2)(2.38*10^-2)^2 = 6.74*10^-6

2) Fe(OH)2 dissolves in water to a small extent according to the equation

Fe(OH)2 = Fe^2+ + 2OH-

The concentration of Fe2+ = weight/mol.wt = 1.15*10^-3/89.86 = 1.28*10^-5 M

if the concentration of Fe2+ is x, then concentration of OH- is 2x i.e., 2.56*10^-5M

therefore, Ksp = (Fe2+)(OH-)^2 = (1.28*10^-5)(2.56*10^-5)^2 = 8.38*10^-15